Engineering Notes on September 2000 IOTA/FLUOR CfA Run
John Monnier

Dates: September 3 - 14, 2000
Observers: John Monnier and Costas Papaliolios

New Memos Generated from Trip
1. Replacing CCDs
2. Installing Anorad Delay Line Cart
3. Removing Anorad Delay Line Cart
4. Planting and Harvesting Anorad Magnets
5. Anorad Screw Table

Summary of the Engineering Performed

1. Replaced the CCD head and chip in FIXED (ironically), because it was not responding to light. We do not know how much worse/better the sensitivity is compared to the new one, but it works (SM/CP/JDM). Conclusion: Either the first replacement CCD we tried is BAD  or the electronic board is bad. We don't know which one.

2. Calibrating Visible Light Piezos around 0 Volts. 0.246 V for 10 HeNe fringes using (visible table) piezo in open loop mode going from -0.006 V to 0.240 volts. If normal incidence, that would be a physical motion of 10*1/4 * 633 nm per 246 mV = 6.4 microns/Volt

3. The "Ozman" light source does not produce visually detectable white light fringes. Must either replace this bubl with something
   else, or use mercury lamp, which required OPD right to within ~2 cm.

4. Autocal white light fringes for S15N15 A delayed:
   Optimum solution: for S15N15 A delayed.  OFFSET: -1.1cm is optimal using Config file with:
     Int. OPD at home, A ... = -104.6
     Int. OPD at home, B ... = 203.20
     ---- these are the numbers in Traub's Jan2000 memo.
   These values must be modified now because we have moved the HOME flag, see end of memo

5. If an anorad magnet falls off its mounting plate, it can be re-affixed with Krazy Glue (but be careful affixing, since magnetic forces
    can make the placement tricky).

6. Tried out the first attempt at a refurbished air pad, using deeper wells around the air ports.  The pad vibrated too much, and needs to
   be to re-worked.  We also tested without the stubs: no difference.

7. Potentiometer R3  on the ELTROL card can be used to tune up the feedback with some effect.

8. Re-wired the polarity of one of the Hall sensors and the order of the motor windings. Replaced 3 bad chips on Harvard Velocity Card
   #1. The second pair of ELTROL + Harvard Velocity Cards now work with both  the old and new delay line cart.
  We used an HC14 chip instead of an F14 chip right after the line receivers -- seems to work fine

9. Pull-down Magnet Measurements for New Table:
   Plate thickness: 1 cm
   Magnet thickness: 3.5 mm
   Length: 5 1/4 inches

   We shoud make one with the plate being 1/8 in shorter, and order some magnets (COS wrote down some #s on the current magnets).

10. We need to order anorad SCREWS.  Lengths: 1 1/8", 1", and 7/8" with 8/32 threads. Much time as spent grinding down screws and
      popping magnets off plates.

11. We've installed the new delay line table on the track, with the damping plates in place. We have not been able to test how well
      the servo loop performs with the damping plates, but may help with  quency vacuum oscillations. The plates clear the tight clearances
      in the U-groove fine, it seems.

12. One of the side magnets for the new table may be lifting off its epoxy. We ended up using one of the (unused) side magnets from
     the old table.

13. We've shortened the bar on the old delay line, which should slightly extend the range of motion.

14. We added a new HOME flag, since both carts are on the track now.
     New home is 43.8 cm different than the other one.
     To correct old Delay.Config files, ADD 87.6 cm to the
     line "Int. OPD at home, A beam ..." and "Int. OPD at home, B beam ..."

15. We removed 1/8" metal shims from the pull-down magnets on the old delay line table and it now moves very smoothly with much
      less air flow. We also only installed 2 of the 4 side magnets. In Vacuum, the necessary flow rates are 1.5 and 1.25 (on right-hand
      flow meters) to float pad and quench oscillations.

16. We discovered that Test Port (TP) 14 (referenced to TP 1) on the ELTROL card is a current sensor, revealing how hard the motor
    is working.   According to the schematics, the conversion is 1 V/Amp. We tried this: when not oscillation and moving at 1 cm/s, the
    measurement said ~1 amp.  When we upped the air flow (it was oscillating) the current was .3 amps. So we are using 3X the power
    with the friction necessary for damping the oscillations.

17. Image quality is still bad in at least one of the telescopes. Suggestion is to try to improve image quality by visually peering
      through alignment telescope and only adjusting the secondary tilt. This should compensate for the tight centration tolerances.

Back to Engineering Notes Page